### v1

# Federated nnU-Net for Privacy-Preserving Medical Image Segmentation

Abstract
The nnU-Net framework has played a crucial role in medical image segmentation and has become the gold standard in multitudes of applications targeting different diseases, organs, and modalities. However, so far it has been used primarily in a centralized approach where the data collected from hospitals are stored in one center and used to train the nnU-Net. This centralized approach has various limitations, such as leakage of sensitive patient information and violation of patient privacy. Federated learning is one of the approaches to train a segmentation model in a decentralized manner that helps preserve patient privacy. In this paper, we propose FednnU-Net, a federated learning extension of nnU-Net. We introduce two novel federated learning methods to the nnU-Net framework – Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg) – and experimentally show their consistent performance for breast, cardiac and fetal segmentation using 6 datasets representing samples from 18 institutions. Additionally, to further promote research and deployment of decentralized training in privacy-constrained institutions, we make our plug-n-play framework public. The source-code is available at https://github.com/faildeny/FednnUNet.
data privacy, federated learning, image segmentation, nnU-Net, nnUNet
1.  Introduction
Medical image segmentation plays a key role in clinical workflows to accurately delineate anatomical structures and help with diagnosis, treatment planning, and disease monitoring [1, 2, 3, 4, 5, 6]. Deep learning methods, especially convolutional neural networks (CNNs), have shown great success in this field. One of the most effective approaches is the no-new-U-Net framework (nnU-Net) [7] which, over the years, has established itself as the go-to framework for medical image segmentation task, evidenced by its multiple first-place finishes on public leaderboards and its widespread adoption as the foundation for further model development across diverse segmentation challenges [8, 9]. nnU-Net extends the widely adopted U-Net architecture [10], which utilizes an encoder-decoder structure with skip connections to capture both semantic and localization information. Going beyond the original U-Net, nnU-Net automatically configures network architectures tailored to a given dataset, provides a robust training schedule and data augmentation strategy, as well as postprocessing methods to further improve the final performance. This “no-new” principle eliminates much of the manual effort associated with hyperparameter tuning and architectural engineering, rendering nnU-Net both powerful and highly convenient across a broad range of biomedical tasks [7, 11, 12, 13]. Despite these advancements, the development of robust segmentation models often requires aggregating extensive datasets from multiple institutions to capture diverse patient populations and imaging protocols. However, the consolidation of such data is frequently impeded by stringent data privacy regulations, ethical considerations, and logistical challenges associated with data sharing. Federated learning (FL) [14] has emerged as a promising solution to these challenges by facilitating collaborative model training across multiple institutions without the need to exchange raw data. Federated learning, however, presents significant challenges when data are heterogeneous across institutions. Traditional FL approaches typically assume that all participants train identical models on similarly distributed data, yet in real-world scenarios, clients often have different computational resources and dataset characteristics. To address this disparity, methods such as HeteroFL [15] have been introduced to enable federated training among clients with varying model capacities, while more recent efforts [16] have focused on mitigating label noise and inconsistencies inherent in federated settings. Collectively, these approaches underscore the growing need for robust FL techniques that can simultaneously manage model and data heterogeneity. Regardless of the significant federated developments in medical domain [17, 18, 19, 20, 21, 22, 23], those methods have not been explicitly tailored for medical image segmentation tasks using nnU-Net. Moreover, nnU-Net’s plug-and-play design, which has made it the gold standard in centralized settings, no longer functions seamlessly in federated environments, necessitating tailored adaptations to overcome the challenges of decentralized, heterogeneous data. To bridge the gap between the decentralized learning paradigms and the robust capabilities of nnU-Net framework, we introduce FednnU-Net, federated extension of nnU-Net that enables collaborative training while preserving data privacy. To the best of our knowledge, it is the first fully-federated and privacy-preserving implementation of nnU-net. We propose two novel methods–Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg). These methods address the challenges posed by variations in dataset properties across different centers, ensuring that each institution maximizes the benefit from shared knowledge training. Our experiments demonstrate that both methods consistently outperform local, non-collaborative training setups for 2D and 3D segmentation tasks across six public and non-public datasets. By extending nnU-Net’s automated pipeline to federated learning and introducing methods to handle data heterogeneity, this work aims to advance the development of robust, scalable, and privacy-preserving medical image segmentation solutions.
2.  Methods
We propose two novel methods to federate nnU-Net’s pipeline for image segmentation: FFE and AsymFedAvg. These two approaches address the challenges of decentralized, heterogeneous data in model training while maintaining nnU-Net’s self-configuring advantages.
2.1.  nnU-Net vs FednnU-Net
nnU-Net is a multi-stage segmentation pipeline that begins with a preparation stage, where dataset fingerprint extraction is performed and a customized training plan (including a tailored network architecture) is generated. This is followed by a training stage and concludes with a post-training stage that selects the best-performing models and optimal image post-processing techniques. To extend nnU-Net into a federated learning setting, we integrate our proposed methods at key points in this pipeline. The first method, FFE, is incorporated into the preparation stage to ensure that data characteristics from all participating institutions are captured and considered when generating the training plan. The second method, AsymFedAvg, operates during the training stage and facilitates the aggregation of model weight updates across nodes with heterogeneous architectures. The following sections provide detailed descriptions of each method. For clarity, we use the terms ”node” and ”client” interchangeably to refer to each institution participating in the decentralized training process, while ”server” describes the coordinating agent responsible for communication and aggregation of inputs from the nodes.
2.2.  Federated Fingerprint Extraction
Figure 1: The Federated Fingerprint Extraction (FFE) process. Each node computes a local fingerprint from its dataset and shares it with the server, which then aggregates them into a global fingerprint. Finally, the server redistributes this global fingerprint to all nodes for local nnU-Net configuration. This process happens once at the beginning of the training stage.
The proposed FFE method enables the analysis of the federated datasets and the configuration of nnU-Net, allowing to determine a unified, one-fits-all training strategy for all data centers in the federated setup. This approach effectively approximates the configuration behaviour in a centralized environment. The dataset’s fingerprint is a crucial component of the nnU-Net’s adaptive configuration, summarizing the key dataset characteristics that are used to determine the model configuration and preprocessing strategies. The FFE approach comprises two key steps:
  • Local Fingerprint Extraction: Each participating node generates their dataset fingerprint by invoking the extraction process of nnU-Net. This local fingerprint encloses the local dataset’s pixel intensity-based metrics, as well as spatial characteristics (such as spacing and voxel size) (see Table 1).
  • Global Fingerprint Aggregation:
    The nodes share their local fingerprint with the server, where these get aggregated to form a global fingerprint. Explicitly, the characteristics are aggregated as follows:
    $\hat{F}={\hat{f}k=g_{k}({f_{k}^{i}}{i=1}^{N})|k\in\mathcal{K}}$ (1)
    where $\hat{F}$ denotes the aggregated fingerprint, $f_{k}^{i}$ represents the $k-th$ fingerprint key from node $i$, and $g_{k}$ is the corresponding aggregation function. Table 1 lists the corresponding $g_{k}$ functions for each fingerprint key. The weighted_mean is calculated with weights proportional to data samples quantity from each participating node.
Table 1: Dataset fingerprint characteristics and aggregation Functions.
Fingerprint Key Aggregation Function $g_{k}$
Pixel intensity
max global max
min global min
mean weighted_mean
median weighted_mean
std weighted_mean
percentile_00_5 weighted_mean
percentile_99_5 weighted_mean
Shape
median_relative_size_after_cropping weighted_mean
shapes_after_crop concatenate
spacings concatenate
Fig. 1 showcases an overview of the FFE process pipeline. As a final step, each node calculates its own nnU-Net training plan based on the received global fingerprint and local hardware characteristics. This process ensures that all nodes will be optimizing either highly compatible or, in most cases, identical network architectures (if the default nnU-Net’s memory constraint has not been altered between nodes).
2.3.  Asymmetric Federated Averaging
In scenarios where nnU-Net is deployed across federated nodes, variations in model architecture can arise due to differences in data characteristics or computational capabilities. To address this inherent heterogeneity, we extend the classical FedAvg approach by developing an Asymmetric Federated Averaging strategy (AsymFedAvg). This method enables nodes to share only the parts of the model that have a common space with the rest of the federation, effectively eliminating the need to constrain all nodes to a single, identical architecture.
At each participating node, we initialize nnU-Net’s adaptive configuration pipeline to generate node-specific U-Net architectures based on the locally extracted dataset fingerprints of the centers. A set of example configurations obtained for each data center is presented in Table 2.
Table 2: Default nnU-Net training configurations for each center in MAMA-MIA dataset
Parameter ISPY1 ISPY2 DUKE NACT
2D Configuration
Batch Size 49 32 12 49
Patch Size [256,256] [320,320] [512,512] [256,256]
Spacing [0.78,0.78] [0.68,0.68] [0.70,0.70] [0.70,0.70]
U-Net Stages 7 7 8 7
Features/Stage [32-512] [32-512] [32-512] [32-512]
Let $\mathcal{K}={1,\dots,K}$ be the set of participating nodes in the federation. For each node $k\in\mathcal{K}$ we define a local model state dictionary as:
$S^{k}={(l_{i}^{k},\theta_{i}^{k})|i\in\mathcal{L}^{k}}$ (2)
where $l_{i}^{k}$ represents the layer identifier, $\theta_{i}^{k}$ denotes the corresponding parameters for node $k$, and $\mathcal{L}^{k}$ is the set of layers in node $k$’s model.
We define a layer matching function $\mathcal{C}$ which identifies the common layers of the models participating in the federation as:
$\small\mathcal{C}({S^{k}}_{k=1}^{K})={l|\forall k,j\in\mathcal{K}:l\in\mathcal{L}^{k}\cap\mathcal{L}^{j}\land\text{dim}(\theta_{l}^{k})=\text{dim}(\theta_{l}^{j})}$ (3)
For each compatible layer $l\in\mathcal{C}({S^{k}}_{k=1}^{K})$, the aggregated parameters $\hat{\theta}_{l}$ are computed as:
$\hat{\theta}_{l}=\frac{1}{|\mathcal{K}l|}\sum{k\in\mathcal{K}_{l}}\theta_{l}^{k}$ (4)
where $\mathcal{K}_{l}$ is the set of nodes that contain layers $l$.
The aggregated model state dictionary $\hat{S}$ is then constructed as:
$\hat{S}={(l,\hat{\theta}l)|l\in\mathcal{C}({S^{k}}{k=1}^{K})}$ (5)
During each federation round $t$, the update process goes as follows:
Algorithm 1 Asymmetric Model Aggregation
1:Each node $k$ sends its state dictionary $S^{k}_{t}$ to the server
2:The server identifies compatible layers using the matching function $\mathcal{C}$
3:The server computes $\hat{S}_{t}$ using the averaging procedure
4:for each node $k$ do
5:     for each layer $l$ in model do
6:         if $l\in\mathcal{C}(\{S^{k}\}_{k=1}^{K})$ then
7:              $\theta_{l,t+1}^{k}\leftarrow\hat{\theta}_{l,t}$
8:         else
9:              $\theta_{l,t+1}^{k}\leftarrow\theta_{l,t}^{k}$
10:         end if
11:     end for
12:end for
The AsymFedAvg allows nodes to maintain their unique architectural components while sharing knowledge via their common space (compatible layers), handling the diverse nature of medical imaging data across different centers. Fig. 2 depicts an overview of the AsymFedAvg pipeline and the process of updating the heterogeneous architectures in each training round.
Figure 2: The Asymmetric Federated Averaging (AsymFedAvg) Pipeline process. The nodes of the federated setup generate a local nnU-Net from the nnU-Net pipeline. During training, the server receives the local nnU-Net weights of each node and aggregates them utilizing the Aggregation Function of AsymFedAvg. Finaly, at the end of each training round, the server distributes the aggregated layers of the network back into the nodes. This process happens throughout the whole training.
2.4.  Implementation
2.4.1 Federated Learning Environment
For the federated communication, we select the Flower framework [24], which is an open-source library designed to streamline federated learning tasks on a cluster of machines. Flower’s high-level API eliminates the need for delving into complex and time-consuming processes such as network communication between nodes and model aggregation, as it already supports multiple FL optimization algorithms such as Federated Averaging (FedAvg), FedProx [25] and Federated SGD (FedSGD) [14].
Additionally, Flower’s client-server architecture is well-suited for our multi-center setup, as it allows seamless integration of local models while maintaining centralized coordination on the server.
2.4.2 FednnU-Net
Our framework maintains full functionality of the original nnU-Net while offering a federated extension that allows experimentation in both real and simulated environments. Its modular structure enables seamless integration not only with the original nnU-Net but also with its customized versions, allowing researchers to adapt their traditional centralized training methods to a federated setup.
2.4.3 Distributed Setup
Our experimental framework consists of 4 computational nodes with varying hardware and system specifications (Table 3), resembling real-world clinical scenarios in low-resource countries. In particular, the availability of GPU device with 8GB of memory, is the requirement for each node in order to run the default nnU-Net training setup. The proposed framework was deployed and tested independently across all nodes, although the final model training is done on a single, multi-GPU cluster to simplify experimental control.
Table 3: Hardware specifications of the computational nodes.
Node CPU (Cores/GHz)* GPU (Memory)** OS ***
Node 1 i7-8700K (12/3.20) GTX 1080 (8GB) Ubuntu 22.04
Node 2 i7-8700K (12/3.20) GTX 1080 (8GB) Windows 10
Node 3 i7-9700K (8/3.60) RTX 2080S (8GB) Ubuntu 22.04
Node 4 i7-9700 (8/3.00) RTX 2080S (8GB) Ubuntu 22.04
*GPU: Graphics Processing Unit, **CPU: Central Processing Unit, ***OS: Operating system
3.  Datasets
We evaluate our proposed methods on three distinct clinical applications, using datasets that serve as valuable testbeds due to their multi-institutional and heterogeneous cohorts. The following sections provide a detailed description of the datasets employed in each use case.
(a) MAMA-MIA
(b) M&M
(c) AIMIX
Figure 3: Example images with annotations from each data center used in the study.
3.1.  Breast MRI
3.1.1 MAMA-MIA
The MAMA-MIA dataset [26] is the largest publicly available collection of breast dynamic contrast-enhanced MRI (DCE-MRI) cases with expert primary tumor segmentations, providing an essential resource for advancing and benchmarking AI algorithms in breast cancer imaging. It comprises 1,506 cases acquired between 1995 and 2016 from four cohorts within The Cancer Imaging Archive (TCIA): ISPY-1 (n=171), ISPY-2 (n=980), Duke (n=291), and NACT (n=64). The annotation process focused on delineating primary tumors and non-mass-enhanced lesions within a single breast, using the first post-contrast phase as the reference—a phase typically associated with maximum enhancement of cancerous tissues that offers superior contrast between lesions and surrounding healthy tissue compared to later phases. However, significant imaging differences among these cohorts (Table 4) pose challenges for training robust models. For example, the NACT cohort predominantly features scans with a 2 mm slice thickness and an average of 60 slices per case, whereas Duke’s samples are mostly acquired with 1 mm thickness and include an average of 169 slices per case. Moreover, the centers used different scanner models and field strengths, further contributing to the variability in imaging protocols.
Table 4: Summary of MRI acquisition parameters across the four centers (ISPY1, ISPY2, DUKE, and NACT) in the MAMA-MIA breast cancer dataset [26].
Center Vendor Scanner Model
Slice
Thickness (mm)
Number of
Slices
Field
Strength (T)
Number of
Studies
ISPY1
GE (67.3%)
Siemens (25.7%)
Philips (7.0%)
Signa Genesis (60.2%)
Other (39.8%)
2.4 [1.5, 4.0] 64 [44, 256]
1.5 (100.0%)
3.0 (0.0%)
171
ISPY2
GE (62.3%)
Siemens (25.7%)
Philips (11.9%)
Signa HDxt (54.7%)
Avanto (12.6%)
Other (32.7%)
2.0 [0.8, 3.0] 106 [52, 256]
1.5 (73.0%)
3.0 (27.0%)
980
DUKE
GE (60.5%)
Siemens (39.5%)
Other (55.7%)
Avanto (24.1%)
Signa HDxt (20.3%)
1.1 [1.0, 2.5] 169 [60, 256]
1.5 (46.7%)
3.0 (53.3%)
291
NACT GE (100%) Signa Genesis (100%) 2.0 [2.0, 2.4] 60 [46, 64]
1.5 (100.0%)
3.0 (0.0%)
64
3.1.2 EuCanImage:
Through the European project EuCanImage [^1] https://eucanimage.eu/, we access a non-public breast MRI collection comprising 574 samples from three different countries. The EuCanImage dataset represents a multi-institutional effort to create standardized imaging biobanks for cancer research and serves here as an external test set to validate our methodology on data acquired under protocols that differ markedly from those in MAMA-MIA. Detailed specifications of the acquisition parameters are provided in Table 5. Notably, while MAMA-MIA includes both axial and sagittal acquisitions, the EuCanImage collection consists exclusively of axial slices. In addition, images in EuCanImage are predominantly acquired from scanner models that differ from those represented in MAMA-MIA, and a significantly larger fraction of the scans are captured at 3T field strength. Although the spatial image characteristics are largely similar between the two datasets, these differences in scanner models, field strengths, and imaging orientations lead to distinct scan appearances, potentially challenging the robustness of segmentation algorithms.
Table 5: Summary of MRI acquisition parameters across the three centers (GUMED, KAUNO, and HCB) in the EUCanImage breast MRI dataset.
Center Vendor Scanner Model
Slice
Thickness (mm)
Pixel
Spacing (mm)
Field
Strength (T)
Number of
Studies
GUMED (Poland) Siemens Aera (100%) 1.00–1.49 1.00–1.49 1.5 (100%) 30
KAUNO (Lithuania)
Philips (88.4%)
Siemens (11.6%)
Ingenia (88.4%)
Avanto (11.6%)
1.50–1.99 (84.5%) 0.50–0.99 (96.6%)
1.5 (11.6%)
3.0 (88.4%)
232
HCB (Spain)
GE (69.2%)
Siemens (30.8%)
Signa HDxt (63.8%)
Aera (22.8%)
Others (13.4%)
2.00–2.49 (87.2%) 0.50–0.99 (98.1%)
1.5 (90.7%)
3.0 (9.3%)
312
3.2.  Cardiac MRI
3.2.1 M&Ms
To further validate our approach, we leverage the Multi-center, Multi-vendor & Multi-disease (M&Ms) cardiac MRI datasets. As a part of our study, we curate a merged version of datasets that were used in the M&Ms challenges [27];[28] by excluding the overlapping samples based on anonymized patient IDs provided by each institution. This curation allows to reach the full potential of this data collection without the risk of replicated samples. The final, accumulated set consists of 543 unique, short-axis cardiac MRI scans with expert annotations from 5 distinct institutions. We make this collection publicly available [^2] This data resource will be released upon the acceptance of this article.. The M&Ms datasets are particularly valuable for validation due to their inherent heterogeneity (Table 6), they include images from four different vendors (Siemens, GE, Philips, Canon), various scanner models (from older systems like Symphony to newer ones like Vantage Orian), and different field strengths (1.5T and 3.0T). This multi-center, multi-vendor nature mirrors the real-world variability in clinical settings, making these datasets ideal for testing the generalizability of our approach.
The technical heterogeneity across centers is evident in variations of scanner models, field strengths and different acquisition protocols. This diversity presents similar challenges to those encountered in our MAMA-MIA analysis, particularly in terms of standardization and harmonization of imaging parameters.
Table 6: Average specifications of images acquired in different centers across the merged collection of M&Ms 1 [27] and M&Ms 2 [28] challenge datasets.
Center Vendor Scanner Model
Slice
Thickness (mm)
Number of
Slices
Field
Strength (T)
Number of
Studies
Hospital Vall d’Hebron Siemens Symphony TIM (63.18%) Avanto Fit (15.48%) Other (21.34%) 9.5 13 1.5 (97.91%) 3.0 (2.09%) 239
Clinica Sagrada Familia Philips Achieva 10 13 1.5 98
Universitätsklinikum Hamburg-Eppendorf Philips Achieva 10 11 1.5 51
Hospital Universitari Dexeus GE Signa Excite (75.24%) Signa HDxt (23.81%) Signa Explorer (0.95%) 9.9 12 1.5 (88.57%) 3.0 (11.43%) 105
Hospital Sant Pau Canon Vantage Orian 10 13 1.5 50
3.2.2 ACDC:
The Automated Cardiac Diagnosis Challenge (ACDC) dataset consists of 150 cardiac short-axis MRI examinations acquired at the University Hospital of Dijon (France) [3]. The dataset is evenly distributed across five groups (30 cases each): normal subjects (NOR), patients with myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (RV). This dataset is particularly valuable for validation due to its well-defined pathological diversity and standardized acquisition protocols. It is used for external validation of models trained on M&Ms dataset.
3.3.  Fetal Ultrasound
3.3.1 AIMIX
With a focus on resource-constrained environments, the AIMIX [^3] https://aimix-erc.eu/ dataset was developed to advance inclusive image analysis in prenatal care. This dataset addresses critical challenges, including the scarcity of African medical data, limited computational resources, and a lack of clinical expertise. The dataset comprises pregnancy blind-sweeps 2D ultrasound scans, along with associated patient metadata such as age and ethnicity. The data span multiple international sites, including Algeria, Egypt, Malawi and Spain, reflecting diverse geographic and healthcare settings.
In this study, we focus on abdominal segmentation, a crucial step in prenatal ultrasound analysis. To enable the evaluation on this multi-center dataset, we provide a set of 130 standard-plane images with masks corresponding to the fetal abdomen [^4] This data resource will be released upon the acceptance of this article.. Our data processing workflow involved detecting the optimal abdominal plane from the each blind-sweep and segmenting it using a 4-point ellipse-fitting algorithm [29], utilizing a semi-automatic software by a researcher with more than five years of experience and reviewed by an expert clinician. Table 7 lists the number of samples and the mean resolution for each center.
Table 7: Summary of the AIMIX and ACOUSLIC-AI datasets, including number of samples, mean resolution, nnU-Net patch size, and UNet stages for each dataset.
Center
Mean Resolution
(h x w)
Number of
Studies
Algeria 418x500 25
Egypt 450x400 20
Malawi 400x500 25
Spain 1228x1276 60
SierraLeone (External) 561x743 178
3.3.2 ACOUSLIC-AI Sierra Leone:
For the external validation, we utilize the publicly available samples from ACOUSLIC-AI [^5] https://acouslic-ai.grand-challenge.org/datasets/ Challenge dataset [30]. We use a collection of 178 cases collected in Sierra Leone from three Public Health Units (PHUs). With a resolution of 744×562 pixels, the images from Sierra Leone sit between the low-end devices used in Algeria and Malawi and the high-end scanners employed in Spain. This intermediate imaging quality makes the dataset an ideal candidate for assessing model generalizability across varying image qualities and clinical environments, ensuring robust validation of AI performance.
4.  Experiments
4.1.  Experimental setups
The training process is conducted across four different setups:
  • Single-center Training:  Each node independently trains its own nnU-Net model using only its local dataset. This setup serves as a baseline, reflecting the performance of a node/institution without participation in the federation.
  • Centralized Training:   All data is aggregated into a single dataset, which is used to train a centralized model. This setup represents the theoretical upper bound of performance, though it is typically infeasible in real-world scenarios due to privacy constraints.
  • Federated Fingerprint Training:   Utilizing our FFE approach (2), the local dataset fingerprints from all nodes are aggregated at the central node into one federated fingerprint and distributed back to each node. As a result, each training site obtains identical network architecture during the nnU-Net’s self-configuration process. This approach approximates the nnU-Net’s behavior in the centralized training setup.
  • Asymmetric Aggregation Training:   Implemented using our AsymFedAvg strategy, allowing nodes to maintain their original nnU-Net configured architectures while sharing compatible parts of their networks.
For our experimentation, we maintain nnU-Net’s default configuration parameters as described in [7]. We test the 2D U-Net architecture on breast MRI and fetal ultrasound data and the 3D architecture on a cardiac MRI use case, consistently employing the default training length of 1000 epochs and using the last model checkpoint for evaluation across all federated nodes. Following other recent works on medical image segmentation [7, 31, 32, 33], we evaluate the models with Dice Similarity Coefficient (DSC) and 95% Hausdorff distance (HD95). For internal validation, we report results as the average of a 5-fold cross-validation, while for external testing, we follow the original nnU-Net approach by generating ensemble predictions from all five folds. Moreover, each experiment uses identical data splits—for example, the n-th fold in the centralized setup corresponds directly to the set of n-th folds from centers used in the local setup. By adhering to these protocols, we try to minimize the influence of other factors on the final results, ensuring that any observed performance variations are primarily due to our federated learning adaptations rather than modifications to the base framework.
4.2.  Results
Table 8: Quantitative results for multi-center MAMA-MIA and EuCanImage datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
Method ISPY1 ISPY2 DUKE NACT external GUMED external KAUNO external HCB
Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
Single-center 0.612 17.60 0.739 19.03 0.616 33.56 0.551 23.24 0.579 35.64 0.636 31.72 0.544 46.30
AsymFedAvg 0.634 16.69 0.739 18.15 0.621 28.70 0.620 21.79 0.557 51.43 0.637 32.43 0.551 49.10
FFE 0.644 16.20 0.742 17.48 0.653 39.54 0.660 17.49 0.582 28.83 0.645 33.52 0.565 40.41
Centralized 0.654 17.28 0.750 19.84 0.657 58.99 0.675 17.35 0.582 52.44 0.644 41.17 0.569 57.98
Table 9: Quantitative results for multi-center M&Ms and ACDC datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
Method Hospital Vall d’Hebron Clinica Sagrada Familia Universitätsklinikum Hamburg-Eppendorf Hospital Dexeus Hospital Sant Pau external ACDC
Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
Single-center 0.908 4.27 0.912 3.62 0.902 6.56 0.896 4.31 0.900 4.34 0.887 4.79
AsymFedAvg 0.909 4.20 0.913 3.70 0.907 4.23 0.897 4.29 0.902 4.07 0.889 4.58
FFE 0.911 4.05 0.914 3.67 0.922 3.30 0.904 3.98 0.904 3.87 0.897 3.87
Centralized 0.910 4.20 0.914 3.53 0.923 3.26 0.905 3.87 0.907 3.81 0.898 3.89
Table 10: Quantitative results for multi-center AIMIX and ACOUSLIC-AI datasets. HD95 values reported in pixels due to missing spacing information. Best results for privacy-preserving methods are in bold.
Method Algeria Egypt Malawi Spain external Sierra Leone
Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
Single-center 0.952 11.87 0.931 24.21 0.926 26.83 0.942 29.82 0.831 53.20
AsymFedAvg 0.951 11.16 0.941 18.29 0.939 16.20 0.944 25.42 0.851 43.44
FFE 0.955 10.23 0.954 14.56 0.943 15.64 0.938 47.58 0.809 46.57
Centralized 0.954 10.23 0.953 14.50 0.943 15.55 0.932 55.76 0.804 54.34
(a) MAMA-MIA cross-center validation matrix
(b) M&Ms cross-center validation matrix
(c) AIMIX cross-center validation matrix
Figure 4: Cross-center DSC scores for isolated model training predictions across multiple centers.
4.2.1 Breast MRI
Table 8 presents the segmentation performance for the breast lesion MAMA-MIA dataset. Both proposed methods, FFE and AsymFedAvg, outperform the local training of individual centers in terms of DSC and HD95 across all datasets. Specifically, FFE method achieves an average DSC improvement of 3.2% on ISPY1, 0.3% on ISPY2, 3.7% on DUKE, and a notable 10.9% on NACT. Additionally, FFE maintains DSC scores close to the centralized approach, while surpassing it in terms of HD95, effectively bridging the performance gap in a decentralized framework. Interestingly for DUKE, AsymFedAvg outperforms every other method in terms of HD95. Since the DUKE dataset comprises samples of substantially different spatial characteristics (pixel resolution approximately 2 times higher, see Table 4), the adaptive nature of the asymmetric aggregation approach allows this center to maintain its initial architecture, while benefiting from the other centers. In this case, the HD95 gets reduced in contrast to unified architecture methods (FFE and centralized setting). The cross-center validation matrix in Fig. 4 further exposes the dataset’s heterogeneity, where both ISPY2 and NACT trained models struggle to generalize well for the DUKE dataset, with the only exception of ISPY2, which has image resolution much closer to DUKE. Sample predictions presented in Fig. 5 show the strong improvement of segmentation performance for NACT dataset, which struggles to perform well with its locally trained nnU-Net model. For DUKE, HD95 improvement can be observed for AsymFedAvg model prediction. External validation results for the EuCanImage dataset show that federated training with the FFE method achieves DSC scores on par with the centralized model. While the segmentation performance for the KAUNO center is comparable to the results achieved within MAMA-MIA dataset, demonstrating good generalizability. The GUMED center exhibits a notable decrease in performance, which is most likely attributed to discrepancies in annotation protocols rather than the data itself. An example of such difference is illustrated in Fig. 7, where fine-grained model predictions fail to align with coarse ground truth regions provided by experts.
Figure 5: Sample predictions of each method for Breast Cancer datasets.
4.2.2 Cardiac MRI
For M&Ms, similarly to previous dataset, both proposed federated learning methods outperform the local, non-collaborative training setup across both tested metrics. The only exception being Clinica Sagrada Familia, for which the HD95 is lower in the local setup by a minimal margin. FFE method consistently achieves the highest scores in terms of DSC for both internal and external datasets, while the adaptive AsymFedAvg does not reflect substantial benefits from federated training. Most likely, it is caused by dominantly homogeneous nature of M&Ms dataset, in which all data centers share similar sample characteristics making a common architecture method more suitable. This effect can be observed as well in the cross-center evaluation of locally trained models (see Fig. 4) where each center managed to provide acceptable results for samples from other, unseen institutions. Moreover, it is important to note that the FFE method, despite the additional privacy requirement manages to provide almost identical performance to the centralized training scenario. Sample predictions for cardiac MRI dataset can be observed in Fig. 6. All tested methods provide consistent results for central slices while having more visible disparities in regions close to the heart’s apex (sample predictions from ACDC) and heart’s atria (sample predictions from Vall d’Hebron).
Figure 6: Sample predictions of each method for cardiac MRI datasets.
4.2.3 Fetal Ultrasound
For fetal abdomen segmentation task, considerable differences can be observed between FFE and AsymFedAvg methods; see Table 10. Due to strong data heterogeneity, i.e. discrepancies in capture devices and image resolution between centers (Table 7), the models trained with data from Algeria, Egypt or Malawi do not generalize well for higher resolution samples from Spain; see Fig. 4c. In this case, having a unified training strategy for all centers does not prove to be the optimal solution. Allowing the largest center, i.e. Spain, to keep its original architecture allows to get a more appropriate model to the characteristics of the external center, i.e. Sierra Leone, while still benefiting from the federated training and contribution of other centers. This effect can be observed on sample predictions presented in Fig. 7. The single-center nnU-Net model often struggled to identify the right region for abdomen delineation, which can be especially observed in the external validation samples from Sierra Leone.
Figure 7: Sample predictions of each method for AIMIX dataset.
5.  Discussion
Our experiments across three diverse imaging use cases - breast MRI (MAMA-MIA), cardiac MRI (M&Ms), and fetal ultrasound (AIMIX) - show that FednnU-Net can reliably match, or even surpass conventional centralized strategies. These results add to a growing body of evidence illustrating that distributed model optimization can achieve comparable accuracy while respecting the complexities of real-world clinical collaborations, such as data privacy and varied institution-specific protocols. Moreover, our findings underscore the critical need for methods that effectively handle data heterogeneity during federated model training.
Whilst other, federated state-of-the-art segmentation methods may achieve higher accuracy with highly customized solutions for specific tasks, the primary objective of this work is to provide an easy-to-use, privacy-preserving solution accessible to researchers with varying levels of expertise. In doing so, we prioritize the broader applicability and real-world viability of the framework over achieving the absolute highest segmentation performance.
The increasing emphasis on privacy-preserving, decentralized settings is rapidly becoming the standard in modern AI solutions — especially within healthcare, where patient data confidentiality is paramount. FednnU-Net aligns with this trend by enabling collaborative training without compromising data privacy. However, several aspects require further investigation:
  1. 1.
    Scalability and Client Diversity: Future studies should test FednnU-Net with a larger number of nodes to better mimic real-world deployments, where numerous and geographically dispersed institutions may challenge model convergence and communication.
  2. 2.
    Heterogeneous Computational Resources: Given that clients often have different processing capabilities, it is crucial to evaluate FednnU-Net under varied hardware constraints. Although our framework already supports such a scenario, this dimension was not explored in this current study.
  3. 3.
    Enhanced Privacy Mechanisms: Beyond our current privacy measures, integrating techniques such as differential privacy [34, 35] and secure multiparty computation (SMPC) could further reduce the risk of data leakage, even in aggregated updates.
In summary, although more accurate medical segmentation techniques exist, FednnU-Net fills a vital gap by balancing segmentation performance with ease-of-use and privacy preservation. As federated learning continues to mature and becomes an increasingly standard practice in healthcare, addressing these limitations through further evaluation and enhancements will be essential for broader adoption and improved robustness in clinical applications.
6.  Conclusion
In this study, we present FednnU-Net, a federated learning framework for 2D and 3D medical image segmentation that bridges the gap between privacy-sensitive setups and high-performance model training. Our framework introduces two methods: Federated Fingerprint Extraction (FFE) and Asymetric Federated Averaging (AsymFedAvg) that unlock the potential of distributed model training under privacy constraints. FFE enables construction of one, universal model architecture akin to centralized training, ensuring robustness across data centers. Alternatively, AsymFedAvg allows tailored model customization for each center to better align with local data characteristics. Our results demonstrate that both FFE and AsymFedAvg offer distinct performance benefits depending on the multi-center dataset characteristics, either achieving performance close to the centralized setting or surpassing it, particularly in highly diverse datasets. On top of that, both approaches preserve nnU-Net’s hallmark self-configuring capabilities, and due to our modular open-source implementation, they allow researchers to seamlessly translate their nnU-Net-derived centralized segmentation methods to the federated setting. Furthermore, we open source our framework and release data enhancing two public dataset collections to allow reproducibility of results. We hope that our contributions provide a robust foundation for further development and evaluation of medical image segmentation models in privacy-constrained settings.